OpenStudy (anonymous):

Solve A = 1/2 p + t for p in terms of A and t

OpenStudy (anonymous):

p=2*(A-t)

OpenStudy (anonymous):

We want to rewrite this where p is the subject. A= 1/2 p + t 1/2p=A-t p=2(A-t)